# Theory and Atomistic Simulation of Electrodeposition

### v1

@@ 1 Introduction
Electrodeposition is the formation of solid structures on the surface of an electrode when an electrochemical potential is applied, and is a viable nanofabrication process alongside more established methods such as nanoimprint lithography 1, 2, pH-driven precipitation 3, 4, and directed assembly 5, 6, 7, 8, 9, 10, 11, 12. Metal electrodeposition is a fundamental electrochemical process, and is an important method with many applications such as carbon dioxide reduction catalysis 13, water splitting 14, fuel cell applications 15, and materials for energy storage and conversion. Metal electrodeposition is also inherently used in various industrial applications, such as electroplating 16, 17, 18, electrowinning 19, 20, 21, and electrocatalysis 22, 23, 24. The key challenge in metal electrodeposition is to control the structure, size, and stability of surface-adsorbed nanostructures on an atomistic scale, which in turn define the reactivity and electrochemical properties of the resulting materials. The establishment of an atomic-scale understanding of the electrodeposition process and dynamics is thus crucial to enable the controlled fabrication of metal nanostructures.
@@
Experimental techniques to track, characterize and harness metal electrodeposition have vastly improved over recent years 25, 26, 27, 28, 29, 30. The complementary use of microscopy approaches, surface spectroscopy methods, and electrochemical analysis provide unprecedented resolution at the nanoscale and, to a more limited extent, resolution in the time domain 26. Simultaneously significant advancements have been made in molecular simulation capabilities and the electronic structure theory of electrified solid-liquid interfaces 31, 32. Yet a large gap between realistic applications, theoretical understanding of dynamics, and atomistic simulation remains. Challenges that require further understanding include electrode-electrolyte interactions, electronucleation and growth mechanisms, and reaction rates and kinetics. Both theory and experiment face challenges when it comes to bridging this gap and reaching an atomic-level understanding of electrodeposition. Theoretical and computational studies must be able to simulate realistic models capable of replicating experimental conditions, accounting for factors such as the electrochemical potential and surface heterogeneity. On the other hand, model experimental studies should ideally be conducted under well-defined and idealized conditions (e.g. atomically-flat electrode interfaces and well-purified electrolytes) to allow for atomistic simulations and theoretical analyses to be applied 33. The synergy between experiment and simulation has the potential to deeply enrich the field, as modeling methods can be refined once information about atomic structure is attained from experiment, while simulations can be used to make predictions that experiments can validate 33. Given the rapid advancement in simulation methodologies and the increasing importance of metal electrodeposition, it is now timely to comprehensively review the vast range of modeling methods available for atomistic simulations. We hope that this review will guide future efforts in this field and help synergize theory and experiment.
@@
Existing reviews tend to focus on specific aspects of computational electrochemistry, such as electron transfer processes 34, 35, modelling methods 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, solvation and solid-liquid interfaces 47, 48, 49, 50, and the electrochemical double layer 51, 52. A useful collection of computational electrochemistry reviews is presented in Koper et al. 53, while Gamburg and Zangari 54 is a recommended resource that covers the theory and practice of metal electrodeposition. However, existing reviews do not focus on metal electrodeposition in particular, nor do they discuss all the various aspects that need to be considered in atomistic simulations of complex electrochemical processes and environments.
@@
In this review, we summarize the key concepts and aspects of electrodeposition, as well as the various state-of-the-art computational techniques that are available for the atomistic simulation of electrodeposition. Finally, we discuss open questions and challenges in the field.
2 Principles of Electrodeposition
@@
The process of metal electrodeposition can be described in four steps (Figure 1): diffusion of metal cations through the solvent after the application of an electric current; adsorption of metal cations at the cathode surface via electron transfer reactions; migration of metal adatoms along the cathode surface; and the nucleation of larger metal nanostructures on the cathode surface.
2.1 Electrodiffusion of Solvated Ions
There are several mathematical relationships that can be used to calculate useful macroscopic quantities in the context of electrodeposition. The Nernst-Planck equation, shown in Equation (1), is a generalization of Fick’s laws 55 and describes the flux, , of a charged particle (such as a precursor cation), due to diffusive and electrostatic forces.
where is the diffusion coefficient of , is the concentration of , is the charge of , is the Faraday constant, and is the electric potential.
The distribution of charged particles within the system will also have an effect on the electric potential, and this relationship can be formulated via the Poisson equation, as shown in Equation (2).
where is the charge density and is the permittivity of the medium.
The Gouy-Chapman equation, shown in Equation (3), can be used to describe the behavior of ions as they approach a charged surface, such as a cathode.
where is the number of electrons transferred, and is the Faraday constant, is the elementary charge, is the Boltzmann constant, is the temperature, and are the concentrations of cations and anions, respectively.
While Equations (1)–(3) provide many valuable insights into the diffusion of solvated ions during electrodeposition, they possess some limitations. The Nernst-Planck equation assumes a dilute solution and that the local electrostatic field is not affected by the ionic concentration, which is likely to not hold at high ionic concentrations. Furthermore, ion-ion interactions are neglected within the formulation, which can significantly influence the diffusive behavior of solvated ions. Additionally, Equation (1) was originally formulated for only binary solutions and would require modifications in order to be applicable for more complicated and realistic systems comprising many ionic species. Despite these limitations, Equation (1) has been widely used to study the time evolution and transport of charged species 56, 57, 58, 59, 60, 61 in experiments.
The Poisson equation assumes solvated ions are stationary and that their motion is not affected by changes in the electrostatic potential. In addition, the distribution of the charge density, , in Equation (2) is assumed to be continuous. Both of these assumptions will not always be accurate in realistic systems, especially with charged metal cations interacting with solvent molecules. Despite this shortcoming, Equation (2) has been used to elucidate various processes during metal electrodeposition, including electrodiffusion, ionic migration and electroconvection 62, 63, 64, 65.
Finally, the Gouy-Chapman equation possesses two major limitations. Firstly, all ions are assumed to be point charges within the system and have the same size. This assumption can result in inaccuracies with solvated metal cations, where particle size can play a part in the final cationic distribution near the cathode surface. Secondly, Equation (3) assumes that the charge at the cathode surface is uniformly distributed in a single plane, and thus neglects more complex situations such as surface heterogeneity and already-adsorbed species.
The limitations of these approaches can be mitigated and/or complemented by atomistic simulation methods, which can be used to provide detailed insights into dynamics and interactions that are not captured by such macroscopic models.
2.2 Electrosorption of Metal Cations
Faraday’s laws of electrolysis 66, summarized in Equation (4), relate the total electrodeposited metal mass, , with the net charge, , within an electrochemical cell:
where is the molar mass. Equation (4) is widely used to calculate various quantities, such as the amount of electrodeposited metal or the time required to achieve a given amount of electrodeposited metal. Faraday’s laws of electrolysis can be manipulated to describe a range of processes, including electroplating, metal corrosion and electrowinning 67, and have been widely applied to experiments 68, 17. However, Faraday’s law does not take into account the interaction between the metal adsorbate and the underlying surface, and only relates with . It therefore assumes that any electrodeposition that occurs is irreversible and effective.
From a more atomistic point of view, the electrosorption of species onto electrode surfaces is typically described using the same theories that are used to treat adsorption in the gas phase, such as -band theory 69. During the adsorption process, bonding and antibonding hybrid orbitals form between the -band of the surface and the valence orbital of the adsorbate. The strength of the binding interaction is determined by the filling of antibonding states 69, 70, 71. However, unlike gas-phase chemistry, where this filling is determined by the number of electrons, at a metal surface, the filling is governed by the energy of the antibonding states with respect to the Fermi level 70 (the energy level at which the probability of an electronic state being occupied by an electron is 50%). As the energies of antibonding states will lie above those of -states, the energy at which the center of the -state distribution lies above the Fermi level can be used to gauge the strength of the binding interaction 70. The higher the energy of the -states with respect to the Fermi level, the higher in energy the antibonding states are and therefore the stronger the binding interaction 70, 72. However, -band theory only considers adsorbate-adsorbent interactions and does not take into account further interactions with the electrolyte or solvent molecules 72; it is thus too simplistic to fully describe electrosorption dynamics 73, 74.
To accurately model metal electrosorption, one must first consider the fundamental differences between adsorption at solid-gas and solid-liquid interfaces. The key difference, in the context of metal electrodeposition, is that once the electrode is negatively polarized, the hydrogen atoms in water molecules will be attracted to the electrode surface (due to their positive dipoles), and these molecules need to be displaced from the cathode surface before metal cations can adsorb. Adsorption isotherms are commonly used to describe adsorption and desorption processes, and determining which adsorption isotherm to use is dependent on the specific characteristics of the system and what assumptions are made by the isotherm for the intended purpose. The simplest example is the Langmuir adsorption isotherm 75, given by Equation (5):
where is the fractional surface coverage (by metal atoms), is the standard electrochemical free energy of adsorption, is the ideal gas constant, is the temperature, and is the activity of the adsorbate in bulk solution. Equation (5) can be rewritten to take into account the applied electrode potential, :
where is the standard free energy of adsorption and is the standard potential of the adsorption reaction 76. While the Langmuir adsorption isotherm provides a first approximation of electrosorption behavior, there are several inherent assumptions that limit its applicability. Firstly, it does not take into account the displacement of solvent molecules, which is a key component of electrosorption. It also assumes the adsorbent surface to be homogeneous, with all adsorbates being assumed to adsorb as a monolayer. Furthermore, all adsorption sites are assumed to be energetically equivalent and interactions between adjacent adsorbed species are neglected. Several extensions to the Langmuir adsorption isotherm exist to mitigate these assumptions 77, 78, 76. For example, a Flory-Huggins-type isotherm can be used to also describe the displacement of solvent molecules during electrosorption 79, 80, 76, while the Frumkin adsorption isotherm 77 can be used to also account for interactions between adjacent adsorbed species, and the Freundlich adsorption isotherm 78 is more appropriate for describing multilayer adsorption on heterogeneous surfaces.
Adsorption isotherms are widely used in experiments 81, 82, 83, 84, 85 to determine parameters such as the maximum adsorption capacity or the relationship between the quantity of particles in a metal deposit and their concentration in an electrolyte. While powerful tools, adsorption isotherms require prior knowledge of the characteristics of the system in some manner and do not take into account electrosorption on an atomistic scale. Atomistic simulations can, however, be used to address these limitations by including factors that are excluded within isotherms.
2.3 Electron Transfer during Electrodeposition
@@
Electron transfer reactions are the process by which metal cations adsorb onto a given cathode surface. Such reactions are termed to be either inner-sphere or outer-sphere. Inner-sphere electron transfer is when the reaction occurs via a strong electronic interaction, such as a covalent bond, and the reactants become connected by a chemical bridge. In contrast, outer-sphere electron transfer (OS-ET) occurs between two ‘unconnected’ reactants, and the electron has to move from one reactant to the other over some space, typically at least a solvent layer from the cathode surface 86 (in the context of electrodeposition). With transition metals, the mechanism of electron transfer for transition metals can be generally assumed to be outer-sphere as, in solution, transition metal cations can form complex coordination compounds with ligands, such as water molecules. The presence of such ligands can create steric hindrance as well as tightly bind the inner coordination sphere around the metal cation, which makes inner-sphere electron transfer less favorable. However, exceptions can arise depending on factors such as the choice of solvent.
@@
Marcus theory is commonly used to describe the rate of electron transfer reactions 87. For multivalent ions, Marcus theory suggests that the simultaneous transfer of multiple electrons is improbable, which implies that the cathodic deposition of multivalent metal cations should occur in a sequence of one-electron steps 87, 88, 89, with the adsorption reaction being the final one-electron step i.e. for a metal element, M: until the neutral atom is reached (i.e. ), rather than . Marcus theory has since been combined with a Newns-Anderson Hamiltonian 90, 91 (vide infra) to describe electrochemical processes by Schmickler and coworkers 88, 89, and has been used to deduce that small monovalent metal cations, such as Ag+, are able to get close to the electrode surface without losing solvation energy as they fit into the water structure well, unlike larger multivalent cations 88, 92. As the valency of metal cations increases, the potentials of mean force sharply increase on approach to the electrode 88, 89. When divalent cations approach the electrode surface, they shed their secondary solvation shells, causing their free energy to rise. This means a close approach of divalent (and multivalent) cations to the electrode surface is energetically unfavorable 89.
Current density, , is typically used as a quantitative measure of the rate at which electrons are transferred between solvated metal cations and the electrode surface during electrodeposition. Current density is typically used instead of just current as larger electrodes can carry more current. The normalization of current by the surface area of the electrode means processes such as electron transfer can be more easily compared on different electrodes.
The current density associated with a first-order reaction is given by Equation (7).
where is the rate constant. Within general (electro)chemical rate theory and transition state theory, obeys the Arrhenius equation 93, 94 and is given as Equation (8).
where is a potential-dependent pre-exponential factor and is the free energy barrier (sometimes referred to as the activation energy), which depends on the electrode potential. accounts for the attempt frequency in transition state theory and may include other effects such as non-adiabatic corrections, solvent dynamics, and nuclear quantum effects 95. It should be noted that the frequency factor, , is often considered in lieu of ; represents the frequency with which a single reactant passes from the reactant side to the product side, and is related to via Equation (9), where is the effective area of the region where the reaction occurs 94.
The reader is directed to He et al. 94 for a more detailed review on the importance of the pre-exponential factor in electrochemistry.
@@
Depending on the degree of electronic coupling between the metal nanostructure and the electrode, OS-ET is classified to be either adiabatic or non-adiabatic. Identifying the adiabaticity of OS-ET for different adsorbate-electrode pairs is of fundamental importance in order to optimize the efficiency and mechanism of electron transfer 96. In the adiabatic regime, is independent of the electron tunneling probability between the metal nanostructure and the electrode, and the rate of OS-ET is independent of the electrode material, assuming there exists a sufficiently strong electronic interaction between the adsorbate and the electrode 94, 97. Using the early version of Marcus theory 87 and collision theory, the value for the pre-exponential factor was typically set to be . Other (and more elaborate) formulae for have been proposed 98, 99, 100, 101, 102 but they typically give rise to the same values in practice 94.
In the non-adiabatic regime, is proportional to the density of electronic states near the Fermi level, as shown in Equation (10) where is the interaction potential and is the electron density.
Factors that can affect include electronic-vibrational interactions 103, the surrounding environment, and the applied potential 104, 105. For metal electrodes, it is usually enough to use the value at the Fermi level as electrons are mainly exchanged at this level 94. OS-ET reactions on pure metal electrodes are often adiabatic 96, while some doubt regarding their adiabaticity on other electrodes remains 96, 106, 107.
2.4 Electronucleation
@@
Once metal atoms have adsorbed onto an electrode surface, they can start to coalesce to form larger nanostructures in a process known as electronucleation. Individual metal adatoms may migrate along the electrode surface and coalesce to form metastable nanoclusters. Furthermore, larger, crystalline nanostructures can form if smaller structures rearrange and amalgamate together, while closely-spaced nanoclusters can also dissemble and feed atoms into existing nanostructures. In contrast, isolated metal adatoms that are not part of a nanostructure and do not move along the electrode surface might indicate the presence of point defect sites on the substrate surface 108. The on-surface dynamics of electronucleation can thus be extremely rich and complex, and atomistic simulations must be able to account for the various thermodynamic and kinetic effects that play a defining part in the size distribution, growth rate, and the rate-determining steps of surface-adsorbed metal nanostructures.
When it comes to modeling the process of electronucleation, both classical and atomistic theories exist to describe the formation of stable nuclei 26, 109, 110, 111. Classical nucleation theory relies on macroscopic physical quantities that are applicable to sufficiently large clusters such that their size, (number of atoms), can be considered a continuous variable. In this case, the Gibbs free energy of nucleation, , is differentiable 110, 111, which allows the rate of nucleation to be predicted 109 using Equation (11):
where is the nucleation rate, is the number of nucleation sites, is the Zeldovich factor (the probability that a nucleus at the top of the barrier for nucleation will progress to form a new phase rather than dissolve 109), is the rate at which atoms attach to the nucleus, is the Boltzmann constant and is the temperature 109. is derived using the assumption that nuclei near the top of the barrier effectively diffuse along the radial axis. By statistical fluctuations, nuclei can either diffusively grow into a larger nucleus that will eventually form a new phase, or can lose atoms and diminish in size.

### v2

# Challenges in the Theory and Atomistic Simulation of Metal Electrodeposition

@@ Introduction
Electrodeposition is the formation of solid structures on the surface of an electrode when an electrochemical potential is applied, and is a viable nanofabrication process alongside more established methods such as nanoimprint lithography 1, 2, pH-driven precipitation 3, 4, and directed assembly 5, 6, 7, 8. Metal electrodeposition is a fundamental electrochemical process with many applications such as carbon dioxide reduction catalysis 9, water splitting 10, fuel cell applications 11, and materials for energy storage and conversion. In particular, electrodeposition plays a crucial role in battery technologies, where it is used in the controlled deposition of metal electrodes 12. Metal electrodeposition is also inherently used in various industrial applications, such as electroplating 13, 14, 15, electrowinning 16, 17, 18, and electrocatalysis 19, 20, 21. Even beyond metals, electrodeposition is commonly used to fabricate non-metallic materials, such as semiconductors 22, ceramics 23, and polymers 24.
Experimental techniques to track, characterize and harness metal electrodeposition have vastly improved over recent years 25, 26, 27. The complementary use of microscopy approaches, surface spectroscopy methods, and electrochemical analysis provide unprecedented resolution at the nanoscale and, to a more limited extent, resolution in the time domain 26. For example, electron microscopy methods have been a popular choice to study metal electrodeposition due to the high resolution they offer 28, 29, 30, 31, 26. Scanning transmission electron microscopy has been shown to be capable of dynamically visualizing the early stages of electronucleation for metals, with structural resolution on the atomic scale and time resolution defined by the sequential analysis of short electrodeposition runs 26. Transmission electron microscopy has also been used to study the electrodeposition of metals at submicroscopic resolution 28, 29, 32, 30, 31, 33. Several studies also report the use of scanning electron microscopy to investigate metal electrodeposition 34, 30, 35. While liquid-cell transmission electron microscopy has made lots of progress in monitoring dynamic electrochemical systems 36, 37, it has limited resolution due to factors such as electron beam-induced gas bubble formation and electron scattering in the liquid 26; in contrast, ex situ aberration-corrected scanning transmission electron microscopy is not only capable of resolving single atoms but can be used to quantify the number of atoms within a particle 38, 26. In contrast, scanning probe methods generate images of surfaces using a physical probe that scans the sample 39. Both atomic force microscopy 40 and scanning tunneling microscopy 41, 42, 43, 44, 45, 46, 47 are capable of analyzing the influence of current on the electrodeposited structure at submicroscopic resolutions. Scanning electrochemical cell microscopy has also been used to study the initial electronucleation stages and mobility of metals 48, 25, 49, 50, and has gained much attention 51, 52, 53 due to its ability to routinely operate at submicroscopic scales 50, 54, 55. Other techniques such as surface plasmon resonance microscopy 56 and dark-field scattering microscopy 57, 58 also exist and have been used to investigate metal electrodeposition. Their performance, however, is restricted by the very small field of view which increases the difficulty in acquiring quantitative data on electronucleation 56, 57. Wide-field surface plasmon resonance microscopy, however, removes this constraint and allows for the growth of hundreds of nuclei to be tracked simultaneously 56.
@@
Simultaneously significant advancements have been made in molecular simulation capabilities and the electronic structure theory of electrified solid-liquid interfaces 59, 60. Yet a large gap between realistic applications, theoretical understanding of dynamics, and atomistic simulation remains. Challenges that require further understanding include electrode–electrolyte interactions, electronucleation and growth mechanisms, and reaction rates and kinetics. Both theory and experiment face challenges when it comes to bridging this gap and reaching an atomic-level understanding of electrodeposition. Theoretical and computational studies must be able to simulate realistic models capable of replicating experimental conditions, accounting for factors such as the electrochemical potential and surface heterogeneity. On the other hand, model experimental studies should ideally be conducted under well-defined and idealized conditions (e.g. atomically-flat electrode interfaces and well-purified electrolytes) to allow for atomistic simulations and theoretical analyses to be applied 61. The synergy between experiment and simulation has the potential to deeply enrich the field, as modeling methods can be refined once information about atomic structure is attained from experiment, while simulations can be used to make predictions that experiments can validate 61. The rapid advancement in simulation methodologies has led to a flurry of atomistic simulation activity in the context of electrocatalysis, while much less attention has been given to electrodeposition. Considering its significant industrial relevance and the need for more controlled growth procedures, atomistic simulations and theories that establish a holistic picture of mass transport, reactivity, and growth are urgently needed. It is therefore timely to review modelling methods available for atomistic simulations of electrodeposition. We hope that this review will encourage and guide future efforts in this field and synergize theory and experiment.
@@
Existing reviews tend to focus on specific aspects of computational electrochemistry, such as electron transfer processes 62, 63, modeling methods 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, solvation and solid-liquid interfaces 75, 76, 77, 78, and the electrochemical double layer 79, 80. A useful collection of computational electrochemistry reviews is presented in Koper et al. 81. Gamburg and Zangari 82 is a recommended resource that covers the theory and practice of metal electrodeposition, and Lin et al. 83 describes nanoscale phenomena of nucleation and crystal growth in electrodeposition; however, neither of these resources discuss how computational simulations can contribute towards understanding electrodeposition phenomena. Existing reviews either do not focus on metal electrodeposition in particular, or they do not discuss all the relevant elementary processes and aspects that need to be considered by atomistic and continuum simulations.
@@
The key challenge in metal electrodeposition is to control the structure, size, and stability of surface-adsorbed nanostructures on an atomistic scale, which in turn define the reactivity and electrochemical properties of the resulting materials. The establishment of an atomic-scale understanding of the electrodeposition process and dynamics is thus crucial to enable the controlled fabrication of metal nanostructures. In this review, we summarize the key concepts and aspects of electrodeposition, as well as the various state-of-the-art computational techniques that are available for the atomistic simulation of electrochemical conditions and electrodeposition processes. Finally, we discuss open questions and challenges in the field.
Principles of Electrodeposition
@@
The process of metal electrodeposition can be described in four steps (Figure 1): diffusion of metal cations through the solvent due the application of an electric current; electrosorption of metal cations at the cathode surface via electron transfer reactions; migration of metal adatoms along the cathode surface; and the nucleation of larger metal nanostructures on the cathode surface. Electrodeposition can also occur without surface migration, taking place directly at steps or kinks or on adsorbed clusters instead.
Despite the significant progress in experimental techniques for studying metal electrodeposition, these methods alone cannot fully resolve the underlying physical and chemical processes at play, particularly at the atomic scale. While electron and scanning probe microscopy provide invaluable insights into nucleation dynamics, surface morphology, and growth kinetics, they are often limited by spatial and temporal resolution constraints, as well as the difficulty of directly probing charge transfer and solvent effects in operando conditions. Moreover, experimental data typically provide snapshots of the electrodeposition process rather than a complete mechanistic understanding of how ions electrodiffuse, electrosorb, transfer charge, and incorporate into the growing metal phase. To bridge this gap, theoretical modeling and simulation play a crucial role in providing a detailed picture of electrodeposition at multiple length and time scales. Computational methods enable the systematic study of electrodiffusion and electrosorption, electron transfer kinetics, and electronucleation mechanisms, offering predictive power that complements and guides experimental investigations. The following sections outline the key theoretical frameworks used to describe these fundamental processes, highlighting the role of atomistic and continuum-scale simulations in advancing the understanding of electrodeposition.
Electrosorption
To accurately model metal electrosorption, one must first consider the fundamental differences between adsorption at solid-gas and solid-liquid interfaces. The key difference, in the context of metal electrodeposition, is that once the electrode is negatively polarized, the hydrogen atoms in water molecules and other (partially) positively charged species will be attracted to the electrode surface, and these molecules need to be displaced from the cathode surface before metal cations can adsorb. Adsorption isotherms have been widely used in experiments 84, 85, 86, 87, 88 to determine parameters such as the maximum adsorption capacity or the relationship between the quantity of particles in a metal deposit and their concentration in an electrolyte. The simplest example is the Langmuir adsorption isotherm 89, which provides a first approximation of electrosorption behavior, but possesses several inherent assumptions that limit its applicability. It does not take into account the displacement of solvent molecules, assumes the adsorbent surface to be homogeneous, and all adsorption sites are assumed to be energetically equivalent with interactions between adjacent adsorbed species being neglected. Several extensions to the Langmuir adsorption isotherm exist to mitigate these assumptions 90, 91, 92. While powerful tools, adsorption isotherms require some prior knowledge of the characteristics of the system and do not take into account electrosorption on an atomistic scale. Atomistic simulations can, however, be used to address these limitations by including factors that are excluded within isotherms.
The electrosorption of species onto electrode surfaces has typically been described using adsorption theories from gas-surface chemistry, such as -band theory 93. However, in electrodeposition, d-bands play a minor role, particularly for metals where -bands do not contribute to the density of states (DOS) at the Fermi level. Rather than solely depending on the -band position, the interaction strength depends on the frontier DOS and the alignment of adsorbate orbitals with metal states near the Fermi level 94, 95. This concept extends beyond transition metals to -band systems, where partially filled states can similarly influence adsorption. Unlike in gas-surface models, interactions with the electrolyte or solvent molecules significantly affect electrosorption 96, making a frontier DOS approach more relevant for describing electrodeposition dynamics 97, 98.
Electron Transfer
During the electrodeposition process, cations undergo electron transfer by which they are reduced to a charge-neutral state.
@@
Such reactions are considered to be either inner-sphere or outer-sphere. Inner-sphere electron transfer is when the reaction occurs via a strong electronic interaction, such as a covalent bond, and the reactants become connected by a chemical bond. In contrast, outer-sphere electron transfer (OS-ET) occurs between two ‘unconnected’ reactants, and the electron has to move from one reactant to the other through space. This typically occurs at least a solvent layer from the cathode surface 99. The mechanism of electron transfer for transition metals can be generally assumed to be outer-sphere as, in solution, transition metal cations can form complex coordination compounds with ligands, such as water molecules, which makes inner-sphere electron transfer less favorable.
The charge transfer kinetics at the interface of an electrode are often modeled using the Butler-Volmer formalism 100, which for the case of single electron transfer can be expressed as:
where , are the reduction and oxidation reaction rates, respectively, is the rate constant, is the transfer coefficient that represents the transition state position, is the Faraday constant, is the applied overpotential, is the gas constant, and is the temperature 101. The transfer coefficient is generally 0.5 for single-electron processes. In the Butler-Volmer formalism, there is a linear relationship between and , with a gradient of , which is independent of the potential. However, studies have shown that this linearity does not generally hold 102 and that a constant value of that is independent of the potential cannot universally be assumed 103, 104.
Marcus 105 developed a kinetic model to describe outer-sphere homogeneous electron transfer, in which the rate expressions can be expressed as:
where is the free energy change on reduction and is the reorganization energy, which requires reorganization of the nuclear configuration of the reactants and solvent to the product state. For multivalent ions, Marcus theory suggests that the simultaneous transfer of multiple electrons is improbable, which implies that the cathodic deposition of multivalent metal cations should occur in a sequence of one-electron steps 105, 106, 107, with the adsorption reaction being the final one-electron step.
@@
Marcus theory in combination with molecular dynamics simulations has been used to deduce that small monovalent metal cations, such as Ag+, are able to get close to the electrode surface without losing solvation energy as they fit into the water structure well, unlike larger multivalent cations 106, 108. As the valency of metal cations increases, the potential of mean force sharply increases on approach to the electrode 106, 107. When divalent cations approach the electrode surface, they shed their secondary solvation shells, causing their free energy to rise. This means a close approach of divalent (and multivalent) cations to the electrode surface is energetically unfavorable 107.
One of the important predictions from Marcus theory is that the transfer coefficient is dependent on the potential. The potential dependence of can be expressed as:
Potential-dependent transfer coefficients have been observed to agree with experiments 102 and have also been used to explain lithium electrodeposition and stripping data 109. However, Marcus theory was originally developed for homogeneous electron transfer reactions, whereby the two reactants are present in the same phase. However, for metal electrodes, for example, the charge transfer reaction is heterogeneous 110, 104 and the reaction kinetics will therefore be dependent on electrons from different energy levels within the conduction band of the metal electrode 111, 112, 113, 114. Based on seminal contributions by Hush, Dogonadze, Kuznetsov, and Gerischer, 115, 116, 117 quantum mechanical rate theories for heterogeneous electron transfer processes have been developed, including what is commonly called the Marcus-Hush-Chidsey model. This model can be expressed with the following rate equation 101:
where is a potential-dependent pre-exponential factor that accounts for the attempt frequency in transition state theory and may include other effects such as non-adiabatic corrections, solvent dynamics, and nuclear quantum effects 118. The reader is directed to He et al. 119 for a more detailed review on the importance of in electrochemistry. In Equation (5), refers to the electronic energy with respect to the Fermi level and the integral captures all electronic contributions to the activation energy weighted by their Fermi-Dirac population, assuming a constant DOS. Marcus-Hush-Chidsey kinetics have been shown to accurately predict Tafel curve data 109, 101.
In metal electrodeposition, electron transfer is often coupled with ion transfer, leading to a deviation from purely electronic Marcus-Hush-Chidsey kinetics. Coupled ion-electron transfer theory extends conventional electron transfer models by explicitly considering the role of ion solvation, desolvation, and migration near the electrode surface. The rate of electrodeposition is influenced by both the electronic structure of the electrode and the reorganization of the solvent environment around the electrodepositing metal ion. Such coupling can lead to deviations from classical Butler-Volmer kinetics, especially in multivalent systems where strong solvation effects and complex reaction pathways modify the potential energy landscape. The incorporation of coupled ion-electron transfer theory into Marcus-Hush-Chidsey kinetics provides a more complete picture of metal deposition and dissolution, particularly in energy storage applications where multivalent ions play a role. Mass transport limitations can manifest in experimental voltammetry as deviations from ideal Butler-Volmer or Marcus kinetics. In particular, limiting current densities and diffusion layer effects can lead to non-ideal Tafel slopes. Computational approaches, such as atomistic simulation methods combined with continuum solvation models, can provide insights into how solvation modulates the potential energy landscape. However, connecting these insights, microscopic ET kinetics, and experimental voltammetric data remains a frontier challenge.
@@
Depending on the degree of electronic coupling between the metal nanostructure and the electrode, OS-ET is classified to be either adiabatic or non-adiabatic. Identifying the adiabaticity of OS-ET for different adsorbate-electrode pairs is of fundamental importance in order to optimize the efficiency and mechanism of electron transfer 120. In the adiabatic regime, is independent of the electron tunneling probability between the metal nanostructure and the electrode, and the rate of OS-ET is independent of the electrode material, assuming there exists a sufficiently strong electronic interaction between the adsorbate and the electrode 119, 121. In the non-adiabatic regime, is proportional to the DOS near the Fermi level. OS-ET reactions on pure metal electrodes are often adiabatic 120, while some doubt regarding their adiabaticity on other electrodes remains 120, 122, 123.
However, Gileadi 124 suggested that metal deposition is unique in that charge is carried across the interface by the metal ion, not the electron. This therefore represents a physical situation that is distinctly different from OS-ET. Gileadi proposed a mechanism, based on the assumption that ions migrate across the double layer under the influence of the electrostatic field created by the applied overpotential, using as the diagnostic criterion. The transfer coefficient quantifies how effectively an applied overpotential lowers the activation barrier for charge transfer, and thus directly influences the nucleation and growth kinetics of metal electrodeposition. Unlike in pure OS-ET reactions, where the solvent dominates the reorganization energy, metal electrodeposition involves strong coupling between the depositing ion and the electrode surface, often with hybridization and partial charge transfer occurring before full incorporation into the growing metal phase. This can lead to a non-trivial dependence of the transfer coefficient on the overpotential and explains observed deviations from the idealized Marcus picture. In particular, the mechanism proposed by Gileadi 124 can also explain the observation that heterogeneous rate constants for metal deposition are often higher or comparable to those of OS-ET reactions, despite the fact that bonds are broken during metal deposition to shed the solvation shell, while typically no bonds are broken in OS-ET. Cations at the electrode are rapidly stabilized by interacting with the electrons in the electrode leading to hybridization and partial charge transfer, followed by the slower process of gradual shedding of the solvation shell. During this process, electrodeposition and electrodissolution will remain in competition.
Electronucleation
@@
Once metal atoms have adsorbed onto an electrode surface, they can start to coalesce to form larger nanostructures in a process known as electronucleation. Individual metal adatoms in partially charged states may migrate along the electrode surface and coalesce to form metastable nanoclusters. Furthermore, larger, crystalline nanostructures can form if smaller structures rearrange and amalgamate together, while closely-spaced nanoclusters can also disassemble and feed atoms into existing nanostructures. In contrast, isolated metal adatoms that are not part of a nanostructure and do not move along the electrode surface might indicate the presence of point defect sites on the substrate surface 125. The potential-driven on-surface dynamics of electronucleation can thus be extremely rich and complex, and atomistic simulations must be able to account for the various thermodynamic and kinetic effects that play a defining part in the size distribution, growth rate, and the rate-determining steps of surface-adsorbed metal nanostructures.